OpenStudy (anonymous):

is the following system consistent or inconsistent? 8x+y+4=0 y=-8x-4

OpenStudy (anonymous):

Re-write the first as y = m x + b and compare.

OpenStudy (anonymous):

8x+y+4 = 0 ---->y= -8x-4 so y= -8x-4 y= -8x-4 the slope is parallel with the same y-intercept, so they are consistent